Izzy Ice Cream has the following price and cost information: $ 5.00 1.35 Price per 2-scoop sundae Variable cost per sundae: IRequired 1 Required 2 Required 3 Required 4 Determine Izzys break-even point in units and sales dollars. sundaes Break-evenRequired 1 Required 2 Required 3 Required 4 Determine how many sundaes must be sold to generate a profit of $6,000. Target saRequired 1 Required 2 Required 3 Required 4 Calculate Izzys new break-even point for each of the following independent scenaRequired 1 Required 2 Required 3 Required 4 Based on the original information, how many sundaes must Izzy sell to generate a
Answer #1

1) Variable cost per sundaes= Ingredients+Direct labor+Overhead= $1.35+0.45+0.20= $2.00

Contribution margin per sundaes= Sales price per sundaes-Variable cost per sundaes

= $5.00-2.00= $3.00

Break-even units= Fixed cost/Contribution margin per sundaes

= $3000/3= 1000 sundaes

Break-even sales= Break-even units*Sales price

= 1000*$5= $5000

Break-even units 1000 sundaes
Break-even sales $5000

2) Break-even units= (Fixed cost+Target profit)/Contribution margin per sundaes

= ($3000+6000)/3= 3000 sundaes

Target sales 3000 sundaes

3-a) New sales price= $5.00-0.50= $4.50

Contribution margin per sundaes= Sales price per sundaes-Variable cost per sundaes

= $4.50-2.00= $2.50

Break-even units= Fixed cost/Contribution margin per sundaes

= $3000/2.50= 1200 sundaes

b) Net fixed costs= $3000-300= $2700

Break-even units= Fixed cost/Contribution margin per sundaes

= $2700/3= 900 sundaes

c) New variable costs= $2.00+0.50= $2.50

Contribution margin per sundaes= Sales price per sundaes-Variable cost per sundaes

= $5.00-2.50= $2.50

Break-even units= Fixed cost/Contribution margin per sundaes

= $3000/2.50= 1200 sundaes

Break-even Point
3a. Sales price decrease by $0.50 1200 sundaes
3b. Fixed costs decreases by $300 per month 900 sundaes
3c. Variable costs increases by $0.50 per sundae 1200 sundaes

4) New variable costs= $2.00+0.30= $2.30

New sales price= $5.00+0.50= $5.50

Contribution margin per sundaes= Sales price per sundaes-Variable cost per sundaes

= $5.50-2.30= $3.20

Break-even units= Fixed cost/Contribution margin per sundaes

= ($3000+10000)/3.20= 4063 sundaes

Break even units 4063 sundaes